# CrackSCF: Lightweight Cascaded Fusion Network for Robust and Efficient Structural Crack Segmentation

### v4

@@ 4.4 Complexity Analysis
To evaluate our LRDS block’s efficacy, we analyzed its computational and parameter efficiency using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS). FLOPs quantify computational load, reflecting demand during inference. Params measure storage requirements, reflecting trainable and stored parameters. FPS gauges real-time processing, indicating frames processed per second under specific hardware conditions.
Table 3 presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS. Our approach achieves notably low values in FLOPs at 9.26G and Params at 4.79M, surpassing the second-best by 38.01% and 18.81%, respectively. This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices. In FPS, our model achieves the second-best performance, processing 36 frames per second, trailing only behind the speed-focused DTrCNet [7]. This highlights our model’s robust real-time capability, crucial for prompt crack image segmentation in practical applications. Overall, our model combines low parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
@@ 4.5 Ablation Studies
We conducted ablation studies on the TUT dataset to assess the effectiveness of our proposed module.
@@
1) Ablation study of different convolution types: As list in Table 4, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution. Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the simple low-rank approximation convolution, achieving the highest FPS. Compared to original convolution operations, our method reduces FLOPs and Params by 89.70% and 73.19%, respectively, without degrading performance. These results indicate that LRDS effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
@@
2) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, we conducted ablation experiments on different segmentation heads, including PID [28], SegFormer [6], UNet [29], and Hamburger head [30], using the TUT dataset. As list in the Table 5, SCFM achieves the best results in segmentation performance metrics. Particularly in terms of F1 and mIoU, SCFM outperforms the UNet head, which achieved the second-best results, by 4.35% and 3.05%, respectively. This indicates that SCFM effectively enhances the model’s segmentation performance by perceiving and integrating local details and long-range dependencies of crack pixels. Furthermore, SCFM requires low computational resources, with Params and FLOPs being 11.95% and 68.84% lower than those of the second-best results, respectively, and achieves the highest FPS. This demonstrates that our SCFM maintains excellent performance while requiring low computational resources.
@@
3) Ablation study of different attention variants in LDE: To validate that our LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, we compared lightweight MSDA with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA. As shown in the Table 6, our method achieves the best results in segmentation performance metrics. Specifically, compared to the common MSDA, our method improves by 1.51% and 1.19% in F1 and mIoU, respectively. Compared to MHA, MSDA outperforms it in segmentation performance, demonstrating the effectiveness of our use of MSDA. Furthermore, whether for MHA or MSDA, when embedded with LR Linear, the required Params and FLOPs significantly decrease. For MSDA, the Params and FLOPs are reduced by 43.43% and 17.70%, respectively. This demonstrates that our LR Linear can effectively reduce the computational requirements needed by the attention mechanism.
@@ 5 Conclusion
This paper proposes a staircase cascaded fusion crack segmentation network (CrackSCF) for pixel-level crack detection. In CrackSCF, we propose the lightweight convolutional block LRDS to construct the MFE, which dramatically reduces the amount of computation and parametric quantities required by the model and is able to efficiently capture the local details of the crack image, and we propose LDE to more efficiently perceive the pixel irregular long-range dependencies in the crack feature maps, in order to more comprehensively fuse the local patterns with the pixel’s long-range dependencies, we also propose SCFM to fully perceive them to generate higher quality segmentation maps. Furthermore, to comprehensively evaluate our model’s robustness and stability across various complex scenarios, we collected the TUT dataset, which includes eight types of crack scenes. Experimental results show that our method performs best on five public datasets and the TUT dataset, particularly excelling in background noise suppression and fine crack segmentation. The computational load and parameter count are reduced to 9.26G and 4.79M, respectively.
Despite our method achieves the low parameter count and computational load, there is still room for improvement in inference speed. In future work, we will further pursue network lightweighting to reduce computational resource requirements and achieve higher inference speeds, aiming to deploy our model on mobile devices for better real-world applications. Additionally, we plan to continue designing an improved SCFM to better fuse and enhance local information and global context for superior segmentation results.

### v5

@@ 3.4 Complexity Analysis
To evaluate the efficacy of the LRDS block, its computational and parameter efficiency was analyzed using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS). FLOPs quantify computational load, reflecting demand during inference. Params measure storage requirements, reflecting trainable and stored parameters. FPS gauges real-time processing, indicating frames processed per second under specific hardware conditions.
Table 8 presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS. Our approach achieves notably low values in FLOPs (9.26G) and Params (4.79M), surpassing the second-best by 38.01% and 18.81%, respectively. This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices. In FPS, our model achieves the second-best performance, processing 36 frames per second, trailing only behind the speed-focused DTrCNet [35]. This highlights our model’s robust real-time capability, crucial for prompt crack image segmentation in practical applications. Overall, our model combines low parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
@@ 3.5 Ablation Studies
To evaluate the strengths of the proposed module over other modules, ablation experiments were performed on the TUT dataset.
Table 9: Comparison of performance metrics using different convolution blocks. Best results are bolded and second best results are underlined.
Method ODS OIS P R F1 mIoU
Original Conv 0.8145 0.8193 0.8294 0.8398 0.8346 0.8437
DS Conv 0.8117 0.8166 0.8239 0.8364 0.8301 0.8417
Low-Rank Conv 0.8154 0.8220 0.8220 0.8451 0.8334 0.8441
LRDS Conv 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
Table 10: Comparison of computational efficiency. Best results are bolded and second best results are underlined.
Method FLOPs $\downarrow$ Params $\downarrow$ FPS $\uparrow$
Original Conv 90.06G 17.86M 23
DS Conv 23.91G 8.38M 35
Low-Rank Conv 21.38G 4.63M 30
LRDS Conv 9.26G 4.79M 36
@@
1) Ablation study of different convolution types: As shown in Table 9, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution. As shown in Table 10, Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the simple low-rank approximation convolution, achieving the highest FPS. Compared to original convolution operations, our method reduces FLOPs and Params by 89.70% and 73.19%, respectively, without degrading performance. These results indicate that LRDS effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
Table 11: Comparison of segmentation performance using SCFM, Single Concat branch, and Single PAF branch.
Method ODS OIS P R F1 mIoU
Single Concat branch 0.8019 0.8092 0.8228 0.8209 0.8219 0.8359
Single PAF branch 0.8137 0.8202 0.8282 0.8297 0.8293 0.8427
SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473
Table 12: Comparison of computational efficiency. Best results are bolded and second best results are underlined.
Method FLOPs $\downarrow$ Params $\downarrow$ FPS $\uparrow$
Single Concat branch 7.21G 4.71M 42
Single PAF branch 8.24G 4.72M 39
SCFM 9.26G 4.79M 36
2) Ablation study in SCFM using different branch: To demonstrate the effectiveness of SCFM in fusing local detail features and long-range information, ablation experiments were performed on the TUT dataset, As shown in Table 11 and 12. Our method, using Concat and PAF dual-branch feature fusion, achieves the best results across all six segmentation metrics. Using only the PAF branch results in a 1.06% drop in the F1 score. While SCFM slightly increases parameter count, computational load, and inference speed, it significantly improves segmentation performance. These results indicate that our module effectively balances performance and efficiency.
Table 13: Comparison of segmentation performance of SCFM with other segmentation heads. Best results are bolded and second best results are underlined.
SegHead ODS OIS P R F1 mIoU
PID 0.7120 0.7198 0.7921 0.6742 0.7284 0.7777
SegFormer 0.7446 0.7634 0.7533 0.7786 0.7658 0.7983
UNet 0.7836 0.7951 0.7829 0.8248 0.8033 0.8222
Hamburger 0.7447 0.7609 0.7075 0.8328 0.7670 0.7969
SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473
Table 14: Comparison of computational efficiency of SCFM with other segmentation heads. Best results are bolded and second best results are underlined.
SegHead FLOPs $\downarrow$ Params $\downarrow$ FPS $\uparrow$
PID 36.69G 5.44M 21
SegFormer 29.72G 6.20M 23
UNet 40.59G 12.34M 20
Hamburger 168.23G 7.45M 14
SCFM 9.26G 4.79M 36
@@
3) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, ablation experiments were conducted on different segmentation heads, including PID [38], SegFormer [36], UNet [27], and Hamburger head [11], using the TUT dataset. As shown in the Table 13, SCFM achieves the best results in segmentation performance metrics. Particularly in terms of F1 and mIoU, SCFM outperforms the UNet head, which achieved the second-best results, by 4.35% and 3.05%, respectively. This indicates that SCFM effectively enhances the model’s segmentation performance by perceiving and integrating local details and long-range dependencies of crack pixels. Furthermore, as shown in 14, SCFM requires low computational resources, with Params and FLOPs being 11.95% and 68.84% lower than those of the second-best results, respectively, and achieves the highest FPS. This demonstrates that our SCFM maintains excellent performance while requiring low computational resources.
Table 15: Comparison of segmentation performance using our lightweight deformable attention in LDE with other variants. Best results are bolded and second best results are underlined.
Attention ODS OIS P R F1 mIoU
Common MHA 0.7919 0.8009 0.7999 0.8240 0.8118 0.8279
MHA + LR 0.7924 0.8003 0.7988 0.8284 0.8134 0.8286
Common MSDA 0.8058 0.8151 0.8052 0.8467 0.8257 0.8373
MSDA 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473
Table 16: Comparison of computational efficiency with different attention variants. Best results are bolded and second best results are underlined.
Attention FLOPs $\downarrow$ Params $\downarrow$ FPS $\uparrow$
Common MHA 15.98G 5.80M 16
MHA + LR 8.87G 4.77M 14
Common MSDA 16.37G 5.82M 28
MSDA + LR 9.26G 4.79M 36
@@
4) Ablation study of different attention variants in LDE: To validate that the LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, lightweight MSDA was compared with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA. As shown in Table 15 and 16, the proposed method achieves the best results in segmentation performance metrics. Specifically, compared to the common MSDA, the proposed method improves by 1.51% and 1.19% in F1 and mIoU, respectively. Compared to MHA, MSDA outperforms it in segmentation performance, demonstrating the effectiveness of utilizing MSDA. Furthermore, whether for MHA or MSDA, when embedded with LR Linear, the required Params and FLOPs significantly decrease. For MSDA, the Params and FLOPs are reduced by 43.43% and 17.70%, respectively. This demonstrates that the LR Linear can effectively reduce the computational requirements needed by the attention mechanism.
Table 17: Comparison of results of CrackSCF networks trained with different loss ratios. Best results are bolded and second best results are underlined
Dice : BCE ODS OIS P R F1 mIoU
1 : 5 0.8129 0.8203 0.8217 0.8458 0.8336 0.8424
1 : 4 0.8049 0.8145 0.8116 0.8365 0.8239 0.8368
1 : 3 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
1 : 2 0.8146 0.8211 0.8227 0.8435 0.8330 0.8438
1 : 1 0.8136 0.8204 0.8231 0.8436 0.8332 0.8431
Only BCE 0.8003 0.8056 0.8349 0.8004 0.8173 0.8332
Only Dice 0.8107 0.8165 0.8302 0.8265 0.8283 0.8409
5) Analysis of different hyperparameter settings in the loss function: Table 17 shows the results of ablation experiments on different weight proportions of the two sub-losses in the loss function. As shown, a Dice loss to BCE loss ratio of 1:3 yields the highest values in five metrics: ODS, OIS, R, F1, and mIoU. Although the P metric is 0.80% lower than using only BCE loss, the F1 metric improves by 2.58%, indicating a better balance between P and R. Therefore, this ratio is adopted during training. In summary, a Dice loss to BCE loss ratio of 1:3 not only enhances pixel-level performance but also improves the global segmentation of crack areas, achieving superior segmentation performance.
3.6 Limitations and Discussions
Through these experiments, it is evident that SCFM effectively perceives and fuses local details and long-range pixel dependencies. The proposed network excels in various metrics on public datasets with simple scenes and performs exceptionally on the complex TUT dataset. The lightweight LRDS block minimizes parameters and computational load, facilitating deployment on resource-constrained devices like drones and smartphones. Additionally, LDE is efficient in capturing irregular long-distance dependencies between cracked pixels. However, the proposed network presents several limitations, as follows:
Figure 10: Examples of CrackSCF misdetections and misses. The critical regions are marked with blue boxes.
  1. 1.
    Although our method generally achieves good crack segmentation with effective background noise suppression, as shown in the Figure 10, our model occasionally misdetects noise regions that resemble cracks in some images. It also fails to detect watermark occlusions with colors similar to the background. Therefore, it is essential to enhance modules such as MFE and LDE to improve the model’s capability to perceive and extract features from critical noise and occlusion areas, thereby further increasing its robustness and stability in complex scenarios.
  2. 2.
    Despite our method achieves the minimum parameter count and computational load, there is still room for improvement in inference speed. For example, our network achieves a processing speed of 36 frames per second, but the DTrCNet [35], which focuses more on inference speed, attains a faster speed of 47 frames per second. Therefore, it is necessary to further simplify convolution operations or the network’s complexity to achieve faster inference speeds while maintaining a good balance between segmentation performance and model lightweighting.
@@ 4 Conclusion
This paper proposes a staircase cascaded fusion crack segmentation network (CrackSCF) for pixel-level crack detection. Within CrackSCF, the lightweight convolutional block LRDS is introduced to construct the MFE. This block dramatically reduces the computational load and parameter quantity required by the model while efficiently capturing the local details of crack images. Additionally, the LDE is proposed to more efficiently perceive irregular long-range dependencies among pixels in the crack feature maps. To comprehensively fuse local patterns with these long-range dependencies, the SCFM is employed to generate higher-quality segmentation maps. Furthermore, to comprehensively evaluate the model’s robustness and stability across various complex scenarios, the TUT dataset, which includes eight types of crack scenes, was collected. Experimental results show that the proposed method performs best on five public datasets and the TUT dataset, particularly excelling in background noise suppression and fine crack segmentation. The computational load and parameter count are reduced to 9.26G and 4.79M, respectively.
Although the proposed method achieves a low parameter count and computational load, there is still room for improvement in inference speed. In future work, network lightweighting will be further pursued to reduce computational resource requirements and achieve higher inference speeds, with the aim of deploying the model on mobile devices for enhanced real-world applications. Additionally, future efforts will focus on designing an improved SCFM to better fuse and enhance local information and global context for superior segmentation results.